# Turbulent dynamos in a collapsing cloud

### v1

@@ Abstract
The amplification of magnetic fields is crucial for understanding the observed magnetization of stars and galaxies. Turbulent dynamo is the primary mechanism responsible for that but the understanding of its action in a collapsing environment is still rudimentary and relies on limited numerical experiments. We develop an analytical framework and perform numerical simulations to investigate the behavior of small-scale and large-scale dynamos in a collapsing turbulent cloud. This approach is also applicable to expanding environments and facilitates the application of standard dynamo theory to evolving systems. Using a supercomoving formulation of the magnetohydrodynamic (MHD) equations, we demonstrate that dynamo action in a collapsing background leads to a super-exponential growth of magnetic fields in time, significantly faster than the exponential growth seen in stationary turbulence. The enhancement is mainly due to the increasing eddy turnover rate during the collapse, which boosts the instantaneous growth rate of the dynamo. We also show that the final saturated magnetic field strength exceeds the expectation from considerations of pure flux-freezing or energy equipartition with the turbulence, scaling as $B\propto\rho^{5/6}$, where $\rho$ is the cloud density. Apart from establishing a formal framework for the studies of magnetic field evolution in collapsing (or expanding) turbulent plasmas, these findings have significant implications for early star and galaxy formation, suggesting that magnetic fields can be amplified to dynamically relevant strengths much earlier than previously thought.
Introduction
Coherent and random magnetic fields are both ubiquitous in stars and galaxies [1, 2, 3, 4]. The leading paradigm that has the potential to explain the origin of these magnetic fields is the turbulent dynamo theory [5, 6]. In a highly conducting turbulent plasma, the small-scale dynamo (SSD) amplifies magnetic fluctuations on a typical time scale of order the turnover time of a turbulent eddy, whereas the large-scale dynamo (LSD) generates fields coherent on scales larger than the turbulent scale on a longer time scale. In galaxies, the two time scales are of order $10\,{\rm Myr}$ and 5–$10\,{\rm Gyr}$, respectively. In protostellar environments, these estimates are less general but the difference in the time scales remains significant.
@@ Introduction
Moreover, magnetic fields coherent on scales of a few kiloparsecs appear to be present in galactic and protogalactic environments even at high redshifts [7, 8, 9, 10, 11]. Their amplification on the corresponding time scale of a few Gyr requires a special explanation as suggested here.
The early production of magnetic fields can have a significant impact on the formation of gas structures. Magnetic fields can hinder fragmentation, and thereby influence the stellar initial mass function [12, 13, 14]. Sufficiently strong magnetic fields can facilitate the formation of jets and outflows from accretion disks [15, 16] removing a significant fraction of mass and angular momentum and thus affecting the stellar mass spectrum [17]. Magnetic fields can influence the evolution of galaxies [18] and contribute to the feedback processes such as galactic winds and fountains [19, 20].
@@
The formation of stars and galaxies starts with the collapse of a gas cloud driven by gravitational, thermal and other instabilities leading to turbulence [21, 22, 23, 24, 25, 26, 27, 28]. The turbulence can be amplified during the gravitational contraction [29, 30, 31, 32, 33, 34, 35, 36]. This presents an intriguing possibility of the dynamo process receiving a boost from the collapse leading to an accelerated growth of magnetic fields. Previous analytical studies [37, 38, 39, 40] did not account for this enhancement of the kinematic dynamo, assuming the kinematic regime to be much shorter than the free-fall time and instead focused on the nonlinear, steady-state dynamo regime [41, 42, 43, 44]. Also, accelerated growth in numerical simulations can be difficult to detect without the right parameter regime and systematic comparisons across parameter variations.
@@
In this letter, we develop an analytical framework for dynamo action in a collapsing cloud to demonstrate that dynamo action leads to a super-exponential growth of the magnetic field strength. Nonlinear regimes of magnetic field evolution in such environments are explored using numerical simulations.
Collapse of a homogeneous cloud
@@
Consider a turbulent cloud undergoing homologous collapse. The position of the fluid element and its density are given by,
@@
Even in the comoving coordinates, the magnetohydrodynamics (MHD) equations are still intricate because of the contribution from the background collapse to the total velocity of fluid element given by
$\mathbf{u}=\mathbf{v}+H\mathbf{r},\qquad H=\mathrm{d}\ln a/\mathrm{d}t,$ (2)
where $\mathbf{v}$ is the peculiar velocity and $H\mathbf{r}$ is the velocity of the collapsing background. Following Ref. [45], we introduce the so-called supercomoving variables denoted here with tilde,
@@
where $p$ is the pressure and $\mathbf{B}$ is the magnetic field. This results in a simpler structure of the governing equations. Note that we use the supercomoving time $\tilde{t}$ instead of $t$ in the analysis and the definition of $\tilde{\mathbf{v}}$ accounts for the evolution of peculiar velocity as $1/a$ in a homologous collapse [46]. Similarly, $\tilde{\mathbf{B}}$ is defined so that the magnetic field amplification via flux freezing, in proportion to $1/a^{2}$, due to the overall collapse is factored out. For an incompressible comoving flow, the MHD equations in supercomoving coordinates follow as
~$\tilde{\nabla}\cdot\tilde{\mathbf{v}}$ $=0,$ (4)
$\frac{\partial\tilde{\mathbf{v}}}{\partial\tilde{t}}+(\tilde{\mathbf{v}}\cdot\tilde{\nabla})\tilde{\mathbf{v}}$ =~$=-\frac{\tilde{\nabla}\tilde{p}}{\tilde{\rho}}+\tilde{a}\frac{(\tilde{\nabla}\times\tilde{\mathbf{B}})\times\tilde{\mathbf{B}}}{4\pi\tilde{\rho}}+\nu\tilde{\nabla}^{2}\tilde{\mathbf{v}},$ (5)
$\frac{\partial\tilde{\mathbf{B}}}{\partial\tilde{t}}$ =~×$=\tilde{\nabla}\times(\tilde{\mathbf{v}}\times\tilde{\mathbf{B}})+\eta\tilde{\nabla}^{2}\tilde{\mathbf{B}},\quad\tilde{\nabla}\cdot\tilde{\mathbf{B}}=0.$ (6)
The gravitational term is balanced by the acceleration of the collapse and, thus, does not appear in the momentum equation. In physical variables, this system corresponds to an adiabatically collapsing sphere, as evident from the scaling of pressure and density in Eq. (3). These equations have the same form as the MHD equations in a stationary background except for the factor $\tilde{a}(\tilde{t})$ in the Lorentz force, and we note that $\tilde{a}(\tilde{t})\equiv a(t)$. The major advantage of these variables is that the induction equation retains its form. This helps to extend the standard kinematic dynamo theory to a homologous collapsing background.
The form of $a(t)$ is governed by the Friedmann equation
$\frac{\ddot{a}}{a}=-\frac{4\pi G}{3}\rho(t),$ (7)
where the dot denotes the time derivative, which can be integrated with the free-fall initial conditions $a(0)=1$, $\dot{a}(0)=0$:
$\arctan\left(\sqrt{(1-a)/a}\right)+\sqrt{a(1-a)}=kt,$ (8)
where $k=\sqrt{8\pi G\tilde{\rho}/3}$. In the limit $a\to 0$, $t$ tends to the free-fall time, $t_{\rm ff}={\pi}/{(2k)}=\left[{3\pi}/{(32G\tilde{\rho})}\right]^{1/2},$ which is inversely proportional to the initial density of the plasma. This solution also leads to the relation between $t$ and $\tilde{t}$. Taking the differential of Eq. (8), we have $\mathrm{d}\tilde{t}=\mathrm{d}t/a^{2}=-k^{-1}a^{-3/2}(1-a)^{-1/2}\,\mathrm{d}a$, which results in
$\tilde{t}=\frac{2}{k}\sqrt{\frac{1-a}{a}}\equiv f(t),$ (9)
@@
$\tilde{a}(\tilde{t})=\frac{1}{1+{k^{2}\tilde{t}^{2}}/{4}}.$ (10)
@@
Turbulent dynamos in a collapsing background
In a stationary environment, the coefficients of the induction equation are independent of time, so dynamo action results in an exponentially fast amplification of an initial (weak) magnetic field at the expense of the kinetic energy of the plasma [49, 50]. The exponential growth continues until the Lorentz force becomes comparable to the forces driving the plasma flow; after that, the exponential growth saturates. Since the induction equation, Eq. (6) in the supercomoving variables has the same form as in a stationary background, the dynamo action causes an exponential amplification of $\tilde{B}$ in terms of the time variable $\tilde{t}$,
$\tilde{B}$ $\propto\exp{\left(\tilde{\gamma}\tilde{t}\right)}.$ (11)
In a dynamo-passive system, $\tilde{\gamma}<0$ because of the magnetic diffusion, but the dynamo action results in $\tilde{\gamma}>0$. Suitable expressions for the growth rate $\tilde{\gamma}$ of the turbulent dynamos (SSD and LSD) can be found in [51]. Using Eqs (3) and (9) for the relations of the supercomoving magnetic field and time to the physical ones, we obtain the corresponding magnetic field strength in the rest frame,
@@
where the factor $1/a^{2}$ is due to the overall collapse and $\tilde{\gamma}=\text{const}$ for both SSD and LSD [51]. Instead of a constant growth rate in a stationary environment, the dynamo action in a collapsing background produces a growth rate increasing with time because of the factor $f(t)$ in the exponent. We refer to this regime as a super-exponential growth of the magnetic field. By the end of the kinematic regime, the Lorentz force affects the plasma flow and the magnetic field strength reaches a stationary saturation level. The magnetic energy density becomes comparable to the turbulent energy density,
@@
in terms of the rms values of $\tilde{B}$ and $\tilde{\rho}\tilde{v}^{2}$, where the factor $\tilde{a}$ is due to the coefficient of the Lorentz force in Eq. (5). In terms of the physical saturated magnetic field strength, assuming $\tilde{\rho}\tilde{v}^{2}=\text{const}$, this implies
$B\propto a^{-5/2}\propto\rho^{5/6},$ (14)
rather than just $a^{-2}\propto\rho^{2/3}$. An important implication of this scaling is that the saturated magnetic fields can significantly exceed the strength predicted by either flux-freezing alone or energy equipartition with the turbulence characteristic of the dynamo in a stationary background:
$B^{2}/(8\pi)\simeq\tfrac{1}{2}\rho v^{2}=a^{-5}\tfrac{1}{2}\tilde{\rho}\tilde{v}^{2}.$
Kinematic dynamos
To verify and refine the analytical arguments, we solved Eqs (4)–(6) numerically, with the induction equation written in terms of the vector potential.
A random flow at an energy-range scale $\tilde{l}_{0}$ is driven by a random force added on the right-hand side of the momentum equation, Eq. (5). We use the forcing function [52] from the Pencil Code [53], rewritten for the supercomoving coordinates (see [51]). The magnitude of the force driving the random flow is independent of the scale while the forcing wavenumbers $\tilde{\mathbf{k}}$ are random in direction and have magnitudes from the uniform distribution in the ranges $1\leq\tilde{k}\leq 3$ (the average $\tilde{k}_{0}=2\pi/\tilde{l}_{0}=2$) for the SSD simulations and $3\leq\tilde{k}\leq 5$ ($\tilde{k}_{0}=4$) for the LSD. Note that the forcing scale decreases with time in the rest frame [41].
@@
We non-dimensionalize the equations with the forcing scale $\tilde{l}_{0}$, the rms random speed $\tilde{v}_{0}$ (even in simulations without a forced random flow), and $\rho_{0}=\tilde{\rho}_{0}$, the plasma density at $t=\tilde{t}=0$. The corresponding unit time is $\tilde{t}_{0}=\tilde{l}_{0}/\tilde{v}_{0}$ and the unit magnetic field is $\tilde{B}_{0}=(4\pi\tilde{\rho}_{0}\tilde{v}_{0}^{2})^{1/2}$. The corresponding kinetic and magnetic Reynolds numbers, $\text{Re}=\tilde{l}_{0}\tilde{v}_{0}/\nu$ and $R_{\text{m}}=\tilde{l}_{0}\tilde{v}_{0}/\eta$, are defined in the supercomoving variables and thus do not change with time. For illustration, $\tilde{t}_{0}\simeq 10^{8}\,{\rm yr}$ and $\tilde{B}_{0}\simeq 4\times 10^{-7}\,{\rm G}$ for $\tilde{l}_{0}=100\,{\rm pc}$, $\tilde{v}_{0}=1\,\text{km}\,{\rm s}^{-1}$, and $\tilde{\rho}_{0}=10^{-24}\,{\rm g}\,{\rm cm}^{-3},$ parameters often used in the simulations of primordial star formation [37].
@@
The kinetic and magnetic Reynolds numbers used are $\text{Re}=R_{\text{m}}=415$ for the SSD simulations, while $\text{Re}=180$ and $R_{\text{m}}=18$ for the LSD. The smaller $R_{\text{m}}$ in the latter case is chosen to support the LSD but not the SSD (which requires $R_{\text{m}}\gtrsim 100$ [56]), so that we can study the two dynamo mechanisms separately.
@@
To reveal the effect of a collapsing background on dynamo action, we use three simulation runs for each of the SSD and LSD: (i) dynamo action in a stationary environment (referred to as the ‘standard’ dynamo) leading to an exponential magnetic field amplification, (ii) background collapse with $t_{\rm ff}=50$ without any random flow, thus no dynamo, and the magnetic field is only amplified due to magnetic flux freezing weakened by the magnetic diffusion, (iii) collapse with $t_{\rm ff}=50$ and driven random flow, so the magnetic field grows due to both dynamo action and the overall collapse of the plasma. The free-fall time chosen is shorter than the time at which the kinematic regime ends for the standard dynamo, $t_{\text{nl}}$ (which depends on the seed magnetic field strength and the dynamo amplification time scale).
@@
Figure 1: Comparison of magnetic field evolution in SSD for different cases. Without collapse, there is exponential growth due to standard dynamo (dash-dotted). With collapse ($t_{\rm ff}=50$) and without forcing (dotted), there is no dynamo; flux freezing competes with resistive diffusion. With both collapse and forcing the dynamo grows the magnetic field super-exponentially (dashed), which is further enhanced by the collapse (solid). The dashed curve factors out the $1/a^{2}$ growth due to overall collapse and emphasizes on the super-exponential growth due to dynamo.
Figure 2: As Fig. 1 but for the LSD with $\text{Re}=180$ and $R_{\text{m}}=18$.
@@
Figures 1 and 2 show the magnetic field evolution in the SSD and LSD, respectively, in the three scenarios. The collapse-induced compression of the magnetic field slows its resistive decay (dotted curves). When the magnetic diffusivity is relatively small ($R_{\text{m}}=415$, Fig. 1), the compression becomes stronger than the decay at the final, accelerated stages of the collapse and the magnetic field strength slightly increases. However, the decay (which is also super-exponential) dominates at all times when the diffusivity is larger, $R_{\text{m}}=18$ (Fig. 2). But, the magnetic field in a collapsing turbulent flow is both compressed by the overall collapse (approximately as $\propto a^{-2}$) and amplified by dynamo action. The contribution of the dynamo is isolated in $a^{2}B$ (dashed), and $\log(a^{2}B)$ increases faster than $t$ showing its super-exponential growth. This confirms that the instantaneous growth rates of both types of the dynamo, SSD and LSD, increase during the collapse. The solid lines in Figs 1 and 2 show the evolution of the physical magnetic field strength confirming that is indeed strongly super-exponential.
I Nonlinear dynamos
@@
To investigate the nonlinear dynamos numerically, we stop the collapse at a certain time $t_{*}$ well before the singular state is reached. After that, the scale factor remains constant at $a_{*}$. If $t_{*}$ occurs in the kinematic dynamo stage, the magnetic field continues growing exponentially (without acceleration) until the saturation. We also considered cases where the dynamo is already in the nonlinear stage at $t_{*}$. Thus, Eq. (10) is modified as
$\tilde{a}(\tilde{t})=\begin{cases}(1+{k^{2}\tilde{t}^{2}}/{4})^{-1}&\text{if}\quad\tilde{t}\leq\tilde{t}_{*},\\ (1+{k^{2}\tilde{t}_{*}^{2}}/{4})^{-1}&\text{if}\quad\tilde{t}>\tilde{t}_{*}.\end{cases}$ (15)

### v2

@@ Abstract
The amplification of magnetic fields is crucial for understanding the observed magnetization of stars and galaxies. Turbulent dynamo is the primary mechanism responsible for that but the understanding of its action in a collapsing environment is still rudimentary and relies on limited numerical experiments. We develop an analytical framework and perform numerical simulations to investigate the behavior of small-scale and large-scale dynamos in a collapsing turbulent cloud. This approach is also applicable to expanding environments and facilitates the application of standard dynamo theory to evolving systems. Using a supercomoving formulation of the magnetohydrodynamic (MHD) equations, we demonstrate that dynamo action in a collapsing background leads to a super-exponential growth of magnetic fields in time, significantly faster than the exponential growth seen in stationary turbulence. The enhancement is mainly due to the increasing eddy turnover rate during the collapse, which boosts the instantaneous growth rate of the dynamo. We also show that the scaling of final saturated magnetic field strength with density robustly exceeds the expectation from considerations of pure flux-freezing. Apart from establishing a formal framework for studying magnetic field evolution in collapsing (or expanding) turbulent plasmas, these findings suggest that during star and galaxy formation magnetic fields can become dynamically relevant much earlier than previously thought.
[…]
@@
Magnetic fields coherent on scales of a few kiloparsecs appear to be present in galactic and protogalactic environments even at high redshifts [7, 8, 9, 10, 11]. Their amplification on the corresponding time scale of a few Gyr requires a special explanation as suggested here.
The early generation of magnetic fields can influence the stellar mass spectrum by suppressing fragmentation and enabling jets and outflows from accretion disks [12, 13, 14, 15, 16, 17, 18]. They also shape galaxy evolution and drive feedback through winds and fountains [19, 20, 21].
@@
The formation of stars and galaxies starts with the collapse of a gas cloud driven by gravitational, thermal and other instabilities leading to turbulence [22, 23, 24, 25, 26, 27, 28, 29], which can be amplified during the gravitational contraction [30, 31, 32, 33, 34, 35, 36, 37]. This presents an intriguing possibility of the dynamo process receiving a boost from the collapse leading to an accelerated growth of magnetic fields. Previous analytical studies [38, 39, 40, 41] do not explore this possibility. Also, accelerated growth in numerical simulations can be difficult to detect without the right parameter regime and systematic comparisons across parameter variations [42, 43, 44].
@@
In this letter, we develop an analytical framework for dynamo action in a collapsing cloud, demonstrating super-exponential growth of the magnetic field strength. Nonlinear regimes of magnetic field evolution in such environments are explored using numerical simulations.
@@
Collapse of a homogeneous cloud.— Consider a turbulent cloud undergoing homologous collapse. The position of the fluid element and its density are given by,
@@
Even in the comoving coordinates, the magnetohydrodynamics (MHD) equations are still intricate because of the contribution from the background collapse to the total velocity of fluid element given by $\mathbf{u}=\mathbf{v}+H\mathbf{r}$. Here $\mathbf{v}$ is the peculiar velocity and $H\mathbf{r}$ is the velocity of the collapsing background with the Hubble parameter $H=\mathrm{d}\ln a/\mathrm{d}t$. Following [45, 46], we introduce the so-called supercomoving variables denoted here with tilde,
@@
where $p$ is the pressure and $\mathbf{B}$ is the magnetic field. This results in a simpler structure of the governing equations. Note that we use the supercomoving time $\tilde{t}$ instead of $t$ in the analysis and $\tilde{\mathbf{v}}$ accounts for the evolution of peculiar velocity as $1/a$ in a homologous collapse [47]. Similarly, $\tilde{\mathbf{B}}$ is defined so that the magnetic field amplification via flux freezing, in proportion to $1/a^{2}$, due to the overall collapse is factored out. In physical variables, this represents an adiabatically collapsing sphere, evident from pressure and density scaling.
For an incompressible comoving flow, the MHD equations in supercomoving coordinates follow as
$\tilde{\nabla}\cdot\tilde{\mathbf{v}}$ $=0,$ (3)
$\frac{\partial\tilde{\mathbf{v}}}{\partial\tilde{t}}+(\tilde{\mathbf{v}}\cdot\tilde{\nabla})\tilde{\mathbf{v}}$ $=-\frac{\tilde{\nabla}\tilde{p}}{\tilde{\rho}}+\tilde{a}\frac{(\tilde{\nabla}\times\tilde{\mathbf{B}})\times\tilde{\mathbf{B}}}{4\pi\tilde{\rho}}+\nu\tilde{\nabla}^{2}\tilde{\mathbf{v}}$, (4)
$\frac{\partial\tilde{\mathbf{B}}}{\partial\tilde{t}}$ $=\tilde{\nabla}\times(\tilde{\mathbf{v}}\times\tilde{\mathbf{B}})+\eta\tilde{\nabla}^{2}\tilde{\mathbf{B}},\quad\tilde{\nabla}\cdot\tilde{\mathbf{B}}=0.$ (5)
The gravitational term is balanced by the acceleration of the collapse and, thus, does not appear in the momentum equation. This choice of incompressible supercomoving flow is consistent with simulations showing that solenoidal turbulence dominates in collapsing primordial cores, being enhanced by the spin-up of turbulent vortices [30]. These equations have the same form as the MHD equations in a stationary background except for the factor $\tilde{a}(\tilde{t})\equiv a(t)$ in the Lorentz force. The major advantage of these variables is that the induction equation retains its form. This helps to extend the standard kinematic dynamo theory to a homologous collapsing background.
The form of $a(t)$ is governed by the Friedmann equation ${\ddot{a}}/{a}=-4\pi G\rho(t)/3,$ where the dot denotes the time derivative, which can be integrated with the free-fall initial conditions $a(0)=1$, $\dot{a}(0)=0$:
$\arctan\left(\sqrt{(1-a)/a}\right)+\sqrt{a(1-a)}={s}t,$ (6)
where ${s}=\sqrt{8\pi G\tilde{\rho}/3}$. In the limit $a\to 0$, $t$ tends to the free-fall time, $t_{\rm ff}={\pi}/{(2{s})}=\left[{3\pi}/{(32G\tilde{\rho})}\right]^{1/2},$ which is inversely proportional to the initial density of the plasma. Taking the differential of Eq. (6), we have $\mathrm{d}\tilde{t}=\mathrm{d}t/a^{2}=-{s}^{-1}a^{-3/2}(1-a)^{-1/2}\,\mathrm{d}a$, which gives the relation,
$\tilde{t}=\frac{2}{{s}}\sqrt{\frac{1-a}{a}}\equiv f(t),$ (7)
@@
$\tilde{a}(\tilde{t})=\frac{1}{1+{{s}^{2}\tilde{t}^{2}}/{4}}.$ (8)
@@
Turbulent dynamos in a collapsing background.—In a stationary environment, the coefficients of the induction equation are independent of time, so dynamo action results in an exponentially fast amplification of an initial (weak) magnetic field at the expense of the kinetic energy of the plasma [50, 51]. The exponential growth continues until the Lorentz force becomes comparable to the forces driving the plasma flow; after that, the exponential growth saturates. Since the induction equation, Eq. (5) in the supercomoving variables has the same form as in a stationary background, the dynamo action causes an exponential amplification of $\tilde{B}$ in terms of the time variable $\tilde{t}$, $\tilde{B}\propto\exp{\left(\tilde{\gamma}\tilde{t}\right)}$. In a dynamo-passive system (when the flow does not support dynamo action), $\tilde{\gamma}<0$ because of the magnetic diffusion, but the dynamo action results in $\tilde{\gamma}>0$. Suitable expressions for the growth rate $\tilde{\gamma}$ of the turbulent dynamos (SSD and LSD) can be found in [52, 53, 54]. Using Eqs (2) and (7) we obtain the corresponding magnetic field strength in the rest frame,
@@
where the factor $1/a^{2}$ is due to the overall collapse and $\tilde{\gamma}=\text{const}$ for both SSD and LSD. Instead of a constant growth rate in a stationary environment, the dynamo action in a collapsing background produces a growth rate increasing with time because of the factor $f(t)$ in the exponent. We refer to this regime as a super-exponential growth of the magnetic field. This is a robust regime which persists even if the supercomoving turbulence in a collapsing cloud is decaying [52].
By the end of the kinematic regime, the Lorentz force affects the plasma flow and the magnetic field strength reaches a stationary saturation level. The magnetic energy density becomes comparable to the turbulent energy density,
@@
in terms of the rms values of $\tilde{B}$ and $\tilde{\rho}\tilde{v}^{2}$, where the factor $\tilde{a}$ is due to the coefficient of the Lorentz force in Eq. (4). In terms of the physical saturated magnetic field strength, assuming $\tilde{\rho}\tilde{v}^{2}=\text{const}$, this implies (see [52] for the decaying case)
$B\propto a^{-5/2}\propto\rho^{5/6}.$ (11)
This magnetic field strength significantly exceeds that due to pure flux freezing alone which leads to $B\propto\rho^{2/3}$, and/or that from considering dynamo in a stationary background.
Kinematic dynamos.—To verify and refine the analytical arguments, we solve Eqs (3)–(5) and (8) numerically, with the induction equation written in terms of the vector potential.
A random flow at an energy-range scale $\tilde{l}_{0}$ is driven by a random force added on the right-hand side of the momentum equation, Eq. (4). We use the forcing function [55] from the Pencil Code [56], rewritten for the supercomoving coordinates (see [52]). The magnitude of the force driving the random flow is independent of the scale while the forcing wavenumbers $\tilde{\mathbf{k}}$ are random in direction and have magnitudes from the uniform distribution in the ranges $1\leq\tilde{k}\leq 3$ (the average $\tilde{k}_{0}=2\pi/\tilde{l}_{0}=2$) for the SSD simulations and $3\leq\tilde{k}\leq 5$ ($\tilde{k}_{0}=4$) for the LSD. We note that the amplitude of the forcing increases and the forcing scale decreases with time in the physical variables, which also allows for the additional driving by the gravitational contraction [26].
@@
We non-dimensionalize the equations with the forcing scale $\tilde{l}_{0}$, the rms speed $\tilde{v}_{0}$ (even in simulations without forcing), and density $\rho_{0}=\tilde{\rho}_{0}$, at $t=\tilde{t}=0$. The unit time is $\tilde{t}_{0}=\tilde{l}_{0}/\tilde{v}_{0}$ and the unit magnetic field is $\tilde{B}_{0}=(4\pi\tilde{\rho}_{0}\tilde{v}_{0}^{2})^{1/2}$. The kinetic and magnetic Reynolds numbers, $\text{Re}=\tilde{l}_{0}\tilde{v}_{0}/\nu$ and $R_{\text{m}}=\tilde{l}_{0}\tilde{v}_{0}/\eta$ are defined in the supercomoving variables and remain constant. For illustration, $\tilde{t}_{0}\simeq 10^{8}\,{\rm yr}$ and $\tilde{B}_{0}\simeq 4\times 10^{-7}\,{\rm G}$ for $\tilde{l}_{0}=100\,{\rm pc}$, $\tilde{v}_{0}=1\,\text{km}\,{\rm s}^{-1}$, and $\tilde{\rho}_{0}=10^{-24}\,{\rm g}\,{\rm cm}^{-3},$ parameters often used in the simulations of primordial star formation [38].
@@
The kinetic and magnetic Reynolds numbers used are $\text{Re}=R_{\text{m}}=415$ for the SSD simulations, while $\text{Re}=180$ and $R_{\text{m}}=18$ for the LSD. The smaller $R_{\text{m}}$ in the latter case support the LSD but not the SSD (which requires $R_{\text{m}}\gtrsim 100$ [59]), so that we can study the two dynamo mechanisms separately.
@@
To reveal the effect of a collapsing background on dynamo action, we use three simulation runs for each of the SSD and LSD: (i) dynamo action in a stationary environment (referred to as the ‘standard’ dynamo) leading to an exponential magnetic field amplification, (ii) background collapse with $t_{\rm ff}=50$ without any random flow, thus no dynamo, and the magnetic field is only amplified due to magnetic flux freezing weakened by the magnetic diffusion, (iii) collapse with $t_{\rm ff}=50$ and driven random flow, so the field grows due to both dynamo action and the overall collapse of the cloud. The free-fall time chosen is shorter than the time at which the kinematic regime ends for the standard dynamo, $t_{\text{nl}}$ (which depends on the seed magnetic field strength and the dynamo amplification time scale).
@@
Figure 1: Comparison of magnetic field evolution in SSD with $\text{Re}=R_{\text{m}}=415$ for different cases. Without collapse, there is exponential growth due to standard dynamo (dash-dotted). With collapse ($t_{\rm ff}=50$) and without forcing (dotted), there is no dynamo; flux freezing competes with resistive diffusion. With both collapse and forcing the dynamo grows the magnetic field super-exponentially (dashed), which is further enhanced by the collapse (solid). The dashed curve isolates the super-exponential growth by removing the factor $1/a^{2}$ arising from the overall collapse.
@@
Figure 1 and Fig. 5 in the end matter show the magnetic field evolution in the SSD and LSD, respectively, in the three scenarios. The collapse-induced compression of the magnetic field slows its resistive decay (dotted curves). When the magnetic diffusivity is relatively small ($R_{\text{m}}=415$, Fig. 1), the compression becomes stronger than the decay at the final, accelerated stages of the collapse and the magnetic field strength slightly increases. However, the decay (which is also super-exponential) dominates at all times when the diffusivity is larger, $R_{\text{m}}=18$ (Fig. 5). But, the magnetic field in a collapsing turbulent flow is both compressed by the overall collapse (approximately as $\propto a^{-2}$) and amplified by dynamo action. The contribution of the dynamo is isolated in $a^{2}B$ (dashed), and $\log(a^{2}B)$ increases faster than $\log(B)$ of the standard dynamo (dash-dotted) showing its super-exponential growth. This confirms that the instantaneous growth rates of both SSD and LSD increase during the collapse. The solid lines in Figs 1 and 5 show the evolution of the physical magnetic field strength confirming that is indeed strongly super-exponential.
@@
Nonlinear dynamos.— To investigate the nonlinear dynamos numerically, we stop the collapse at a certain time $t_{*}$ well before the singular state is reached. After that, the scale factor remains constant at $a_{*}$. If $t_{*}$ occurs in the kinematic dynamo stage, the magnetic field continues growing exponentially (without acceleration) until the saturation. We also considered cases where the dynamo is already in the nonlinear stage at $t_{*}$. Thus, Eq. (8) is modified as
[…]